OpenStudy (anonymous):

Unit Conversion What is 5131 feet^2 into kilometers^2

OpenStudy (anonymous):

\[\frac{ 5131ft^2 }{ 1 }\times \frac{ ?km}{ 1ft }\times \frac{ ?km }{ 1ft }\] 1ft=?km You can search online to get the ?(how much km is 1ft) and put it in the conversion. I put ft x ft in the denominator because it would cancel with the ft^2 out. I put km x km on top because you want km^2 Denominator=numerator in a conversion problem except for the first one.